SCI Question 4

SOLVE FOR THE VALUE OF NET INCOME

 

Gross Sales = 9,000,000 pesos

COGS= 120% of Distribution Expenses  

Distribution Expense= 20% of OCI  

Administrative Expense = 5% of Gross Sales  

Income Tax Expense= 25% of Operating Profit

Other Comprehensive Income= 2,000,000 pesos


1
Expert's answer
2021-09-28T16:57:52-0400

Gross Sales = 9,000,000 pesos

COGS= 120% × 400,000 = 480,000 pesos 

Distribution Expense= 20% × 2,000,000 = 400,000 pesos

Administrative Expense = 5% × 9,000,000 = 450,000 pesos

Income Tax Expense= 25% × 7,670,000 = 1917500 pesos


Gross Profit = Gross Sales - COGS

= 9,000,000 - 480,000 = 8,520,000 pesos


Operating Profit =Gross Profit - Operating Expenses

= Gross Profit - Distribution Expense - Administration Expense

= 8,520,000 - 400,000, -450,000 = 7,670,000


Net Income = Gross Profit - all other expenses

= Gross Profit - Distribution Expense - Administration Expense - Income Tax Expense

= 8,520,000 - 400,000 -450,000 -1,917,500 = 5,752,500 pesos

Net Income = 5,752,500 pesos